For the reaction 2CO(g)+O2(g)⇌2CO2(g), the value of K at a certain temperature is 1300.

Calculate [CO2] at equilibrium when [CO] = 9.0×10−2 mol/L and [O2] = 0.527 mol/L
Answer #1

Equilibrium constant K = [CO2]2/ [CO]2 * [O2] = 1300

i.e. [CO2]2 = 1300 * [CO]2 * [O2] = 1300 * (9.0×10−2)2 * 0.527 =5.549

Hence, at equilibrium [CO2] = [5.549]1/2 = 2.356